# NextLevelBERT: Investigating Masked Language Modeling with Higher-Level Representations for Long Documents

### v1

@@ Abstract
While (large) language models have significantly improved over the last years, they still struggle to sensibly process long sequences found, e.g., in books, due to the quadratic scaling of the underlying attention mechanism. To address this, we propose NextLevelBERT, a Masked Language Model operating not on tokens, but on higher-level semantic representations in the form of text embeddings. We pretrain NextLevelBERT to predict the vector representation of entire masked text chunks and evaluate the effectiveness of the resulting document vectors on three task types: 1) Semantic Textual Similarity via zero-shot document embeddings, 2) Long document classification, 3) Multiple-choice question answering. We find that next level Masked Language Modeling is an effective technique to tackle long-document use cases and can outperform much larger embedding models as long as the required level of detail is not too high. We make model and code available.
NextLevelBERT: Investigating Masked Language Modeling with Higher-Level Representations for Long Documents
@@ 1 Introduction
Beyond word-level predictions, the original BERT architecture Devlin et al. (2019) also includes next sentence prediction as one of the pretraining tasks, but this still is based on individual tokens as input. Masked language modeling has previously been explored at the level of short spans of tokens Joshi et al. (2020), but again assuming tokens as individual inputs.
We present NextLevelBERT, a masked language model operating entirely at larger timescales, with inputs that are not individual tokens but rather generic chunks of text (e.g., sentences or larger spans of tokens). For each chunk, we obtain a representation using a pretrained encoding model. Some of these chunks are then masked and the respective representations need to be predicted based on the context. Our results show that applying masked language modeling to higher-level semantic representations results in meaningful document vectors. NextLevelBERT uses only a third of the number of parameters of its strongest contender, but showcases stronger performance than other commonly used approaches during downstream testing as long as the task does not require a high level of detail at the token level.
@@ Discussion
With the typical input length of pretrained Transformer-based language models of 512, as well as our own model’s input length of 512, the overall receptive field of the model can be stretched to a maximum of $512\times 512=$ 262,144 tokens. This may be extended further by adopting well-known techniques for larger context sizes, which NextLevelBERT can then amplify. For instance, if a basic token-level LM’s context size is extended from 512 to 4,096 tokens, this increases NextLevelBERT’s effective context size quadratically to 16,777,216 tokens.
@@ 3.1 Pretraining
We pretrain each version of the model for 20 epochs on the encoded books3 dataset for up to five days on four Tesla V100-SXM2-32GB. Pretraining time is shorter for larger chunk sizes. We use a cosine learning rate schedule (max. $10^{-4}$) with linear warm-up (first 5% of the total number of steps) and the AdamW (Loshchilov and Hutter, 2018) optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay: $0.01$). Unless otherwise stated, we use a global batch size of 512, a masking rate of 15% and a smooth L1 loss. By default, to encode the input chunks into vectors, we use the MiniLM-L6 model from the Sentence Transformers library [^1] https://www.sbert.net/index.html. In practice, we encode the input chunks during preprocessing, since this can be done very efficiently in large batches over multiple GPUs and can be cached for later reuse. We initialize the NextLevelModel’s weights with the pretrained weights of the chunk encoder, since we found that this sped up convergence (see Section 4.5).
@@
— Sentence-based N/A 384 80 30.97 50.88 58.14 73.53 33.37
— 16 N/A 384 80 25.63 43.66 58.14 73.53 32.68
— 32 N/A 384 80 38.71 59.95 58.14 73.53 33.07
— 64 N/A 384 80 50.36 73.18 58.14 73.53 32.68
— 128 N/A 384 80 59.41 80.85 58.14 73.53 32.78
— 256 N/A 384 80 56.81 78.47 58.14 73.53 32.20
— 512 N/A 384 80 29.28 49.00 58.14 73.53 32.98
SBERT: MiniLM 512 384 80 30.94 48.37 58.14 73.53 32.39
Longformer-base 4,096 768 595 00.55 01.33 58.14 73.53 26.07
Nomic Embed 8,192 768 550 65.07 84.90 73.26 79.28 35.28
Table 2: Evaluation results (in %). Best results per dataset and metric are presented in bold. The sentence-based NextLevelBERT model’s context size is estimated from Books3’s average sentence length. For the Chunk Mean approach, there is no overall context limit due to the simple mean-pooling aggregation.
@@ 3.2 Downstream Evaluation
For MuLD-Mov. and QuALITY, we add a simple classifier head with a hidden layer size of 768 and ReLU activation and subsequently fine-tune the entire model with a learning rate of $5\times 10^{-5}$. For QuALITY, we concatenate the document vector with an explicit encoding of the question and candidate answer and predict four scores via softmax on the minibatch-level, one per question-answer pair. For MuLD-Mov., we instead concatenate the encoded question: "Is the character <NAME> behaving like a hero or a villain?", where <NAME> is replaced by the actual character’s name. We use a sigmoid function (threshold at 0.5) to predict the binary label. Due to limited compute resources, we report the results of a single run.
@@ 3.3 Baselines
is a text embedding model for document lengths up to 8,192 tokens. Based on a BERT architecture, the model is trained via Masked Language Modeling and both unsupervised as well as supervised contrastive textual similarity tasks. We choose this model as a highly competitive baseline with a long context window yet relatively small size at 137M parameters.
@@ 4 Results
- MiniLM-256 384 65.48 85.57 76.74 80.77 29.67
- MPNet-256 768 02.36 04.26 58.14 73.53 33.95
- DistilRoBERTa-256 768 43.10 66.73 58.14 73.53 33.27
SBERT: MiniLM 384 30.94 48.37 58.14 73.53 32.39
SBERT: MPNet 768 44.33 67.19 58.14 73.53 32.59
SBERT: DistilRoBERTa 768 27.93 45.87 58.14 73.53 33.85
Table 3: Evaluation results (in %) for NextLevelBERT pretrained on the text chunk embeddings of various encoder models. All next-level models use text chunks of size 256.
@@ 4.1 Downstream Task Evaluation
On the MuLD-Mov. dataset, four out of seven NextLevel models as well as Nomic Embed are able to surpass chance level (58%, i.e., a majority class classifier). All other models fail to do so, even the Avg. SBERT baseline, which does well on the other datasets. Notably, our sentence-based NextLevelBERT also outperforms the Longformer baseline from Hudson and Al Moubayed (2022) (82.58 F1), the original benchmark paper, without any task-specific preprocessing. To attain their score, the authors first extracted the most important passages based on character name frequency. We hypothesize that a good performance on MuLD is dependent on a model’s ability to process the entire text as well as exchange contextual information internally.
On the QuALITY dataset, Nomic Embed is the strongest model, presumably due to its sufficient context size and the fact that it can provide token-level details to answer the questions. The Avg. SBERT baseline reaches a higher accuracy than NextLevelBERT. Since the chunks that are mean-pooled by Avg. SBERT are also the inputs to NextLevelBERT, a lower performance suggests that NextLevelBERT may actively disregard some of the incoming information. This is likely due to the fact that during pretraining, NextLevelBERT learned to focus on reconstructing general events and to ignore local details that would be hard to predict from the chunk-level context. QuALITY consists of questions such as “Why does Joseph lie about the water supply?”, which require non-trivial consideration of local details. In general, this is a difficult dataset for all the tested models and the results are generally barely above the chance level of 25%. This is partly reflected in the dataset’s leaderboard [^4] https://nyu-mll.github.io/quality/, where end-to-end encoder-only models without additional preprocessing or fine-tuning on related datasets report similar accuracy. The dataset’s current SOTA mostly relies on large ($\geq$7B) decoder-only models, training on a large-scale task-specific reading comprehension corpus, and/or retrieval augmentation.
@@ 4.2 Impact of Chunking Strategy
The results in Table 2 illustrate how the optimal choice for dividing the input text into chunks can vary between use cases. We hypothesize that there is a trade-off between increasing the receptive field by using larger chunks and losing information due to higher levels of compression necessary to encode longer text chunks in fixed-size sentence embeddings. This is supported by the fact that in the cases where the next-level models’ context size is not sufficient, performance quickly degrades. A notable exception is NextLevelBERT with sentence-based chunking on the MuLD dataset. With an average context size of $\sim$12k tokens and MuLD’s average document length being 34,706 tokens, it still outperforms all other models. A possible explanation is that the movie scripts that MuLD consists of are dialogue-heavy and the sentence-based chunking plus the contextualization abilities of the model contribute to successfully predicting the role of the character without seeing the whole document.
@@ 4.3 Impact of Encoder Model
Downstream results for NextLevelBERT with different text encoder models are compared in Table 3. We observe that the effectiveness of the original encoder model does not automatically translate to similar outcomes in the next-level setting. Surprisingly, pretraining on MiniLM leads to the best downstream results, although it is the smallest model in terms of parameters as well as embedding dimensionality. A possible explanation is that capturing the intricacies of a 384-dimensional embedding space during pretraining is easier than for those of a 768-dimensional space.
@@ 4.4 Impact of Document Length on Model Performance
To investigate how the models perform with regard to the length of the document, we visualize the MRR on BookSum per document length in Figure 2. Here, 1(a) provides a general plot covering all document lengths, while 1(b) shows a more fine-grained view of documents shorter than 768 tokens to study if the design for long documents comes at the cost of degraded performance on shorter texts. NextLevelBERT outperforms the other models, except Nomic Embed, across all document lengths with a margin of up to 20% MRR. Even on the shorter documents, it generally improves over the other models starting from 341 tokens. On the longest documents, all models exhibit a MRR of 1, but this is due to the low number of samples at these lengths. Notably, Longformer seems unable to generate any useful document embeddings for any sequence length out-of-the-box.
@@ 4.5 Pretraining Effects
To measure the effect of initializing NextLevelBERT with the chunk encoder weights, we train two versions of the same model, one initialized with the weights of the sentence encoder and one with random initialization. We compare their downstream performance in Table 4. For the MuLD-Movie and the QuALITY tasks, the model with random initialization does not score above chance level. On BookSum it is able to reach scores comparable to the weaker baselines in Table 2, which shows that this version can still capture some general document-level semantics despite failing on tasks that require more specific information. Together, these results reveal that initializing NextLevelBERT with the chunk encoder weights is essential to its downstream capabilities.
BookSum MuLD-Mov QuALITY
NextLeveBERT MRR HR@10 Acc F1 Acc
with 15% 65.48 85.57 76.74 80.77 29.67
with 30% 61.14 82.10 63.95 76.34 30.45
with 40% 64.21 84.81 80.23 83.50 30.35
simpler masking 00.77 01.29 58.14 73.53 27.92
random init. 37.60 59.57 58.14 73.53 24.71
Table 4: Results for alternative masking rates (1-3), replacing all selected inputs with the [MASK] vector (4), and a randomly initialized model (5). All models use text chunks of size 256.
@@ 5 Related Work
Numerous approaches have been proposed to increase the context length by making the attention mechanism more scalable, typically by sparsifying or approximating the attention operation: Beltagy et al. (2020) use a combination of local and global attention to circumvent the quadratic scaling. They train the Longformer, a BERT-based LM that can incorporate a context size of up to 4,096 tokens, showing improved results on long document classification. Building on the Longformer, Xiao et al. (2022) pack multiple documents into the same sequence and mask salient sentences to generate multi-document summaries. The Poolingformer (Zhang et al., ) adds a second level of pooling-combined attention to achieve linear computational complexity, while the Reformer (Kitaev et al., 2019) proposes attention with locality-sensitive hashing to reduce the attention operation to linearithmic time. Finally, HyperAttention (Han et al., 2023) efficiently approximates the attention matrix and enables computing attention in near-linear time.
@@ 6 Conclusion
Successfully modeling long-range dependencies is an important step towards building language models with general language understanding. Inspired by the idea of modeling language at a coarser resolution than that of tokens, we have proposed NextLevelBERT, and transferred masked language modeling from the token level to general text representation vectors. We pretrained NextLevelBERT on a large collection of books and evaluated the resulting text representations on three diverse downstream tasks in a long document setting. Our results show that next level masked language modeling leads to sensibly contextualized text chunk representations and generates useful long-document vectors, especially for tasks that do not require specific token-level information. Additionally, NextLevelBERT is very parameter-efficient, using only a third of the parameters of the Nomic Embed model while maintaining similar performance.
Combining the predictions of a next-level model with a token-level component based on the required granularity could be a way of preserving token-level details. Furthermore, pretraining NextLevelBERT on data from more diverse domains could improve its performance in general and extend its usefulness to a wider range of tasks. Lastly, while not straightforward, adapting decoder-only models to the next level could translate similar gains as in the MLM setup to generative language models.

### v2

# NextLevelBERT: Masked Language Modeling with Higher-Level Representations for Long Documents

@@ Abstract
While (large) language models have significantly improved over the last years, they still struggle to sensibly process long sequences found, e.g., in books, due to the quadratic scaling of the underlying attention mechanism. To address this, we propose NextLevelBERT, a Masked Language Model operating not on tokens, but on higher-level semantic representations in the form of text embeddings. We pretrain NextLevelBERT to predict the vector representation of entire masked text chunks and evaluate the effectiveness of the resulting document vectors on three types of tasks: 1) Semantic Textual Similarity via zero-shot document embeddings, 2) Long document classification, 3) Multiple-choice question answering. We find that next-level Masked Language Modeling is an effective technique to tackle long-document use cases and can outperform much larger embedding models as long as the required level of detail of semantic information is not too fine. Our models and code are publicly available online.
NextLevelBERT: Masked Language Modeling with Higher-Level Representations for Long Documents
@@ 1 Introduction
Beyond word-level predictions, the original BERT architecture Devlin et al. (2019) also includes next sentence prediction as one of the pretraining tasks, but this still is based on individual tokens as input. Masked language modeling has previously been explored at the level of short spans of tokens Joshi et al. (2020), but again assuming tokens as inputs.
We present NextLevelBERT, a masked language model operating entirely at larger timescales, with inputs that are not individual tokens but rather generic chunks of text (e.g., sentences or larger spans of tokens). For each chunk, we obtain a representation using a pretrained encoding model. Some of these chunks are then masked and the respective representations need to be predicted based on the context. Our results show that applying masked language modeling to higher-level semantic representations results in meaningful document vectors. NextLevelBERT uses only a third of the number of parameters of its strongest contender, but showcases stronger performance than other commonly used approaches during downstream testing as long as the task does not require a high level of semantic detail at the token level.
@@ Discussion
With the typical input length of pretrained Transformer-based language models of 512, as well as our own model’s input length of 512, the overall receptive field of the model can be stretched to a maximum of $512\times 512=$ 262,144 tokens. Theoretically, this may be extended further by adopting well-known techniques for larger context sizes, which NextLevelBERT can then amplify. For instance, if a basic token-level LM’s context size is extended from 512 to 4,096 tokens, this could increase NextLevelBERT’s effective context size quadratically up to 16,777,216 tokens.
@@ 3.1 Pretraining
We pretrain each version of the model for 20 epochs on the encoded books3 dataset for up to five days on four Tesla V100-SXM2-32GB. Pretraining time is shorter for larger chunk sizes. We use a cosine learning rate schedule (max. $10^{-4}$) with linear warm-up (first 5% of the total number of steps) and the AdamW (Loshchilov and Hutter, 2018) optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay: $0.01$). Unless otherwise stated, we use a global batch size of 512, a masking rate of 15% and a smooth L1 loss. By default, to encode the input chunks into vectors, we use the MiniLM-L6 model from the Sentence Transformers library [^1] https://www.sbert.net/index.html. In practice, we encode the input chunks during preprocessing, since this can be done very efficiently in large batches over multiple GPUs and can be cached for later reuse. We initialize the NextLevelModel’s weights with the pretrained weights of the chunk encoder, since we found that this is crucial for convergence (see Section 4.5).
@@
— Sentence-based N/A 384 80 30.94 50.85 58.14 73.53 32.51
— 16 N/A 384 80 25.63 43.64 58.14 73.53 31.93
— 32 N/A 384 80 38.64 59.93 58.14 73.53 31.69
— 64 N/A 384 80 50.35 73.18 58.14 73.53 31.81
— 128 N/A 384 80 59.37 80.85 58.14 73.53 31.95
— 256 N/A 384 80 56.84 78.49 58.14 73.53 31.89
— 512 N/A 384 80 29.25 48.98 58.14 73.53 31.67
SBERT: MiniLM 512 384 80 30.87 48.35 58.14 73.53 31.96
Longformer-base 4,096 768 595 01.92 04.48 58.14 73.53 26.41
Nomic Embed 8,192 768 550 65.07 84.90 60.47 73.85 36.58
Table 2: Evaluation results (in %, averaged over five fine-tuning runs). Best results per dataset and metric are presented in bold. The sentence-based NextLevelBERT model’s context size is estimated from Books3’s average sentence length. For the Chunk Mean approach, there is no overall context limit due to the simple mean-pooling aggregation.
@@ 3.2 Downstream Evaluation
For MuLD-Mov. and QuALITY, we add a simple classifier head with a hidden layer size of 768 and ReLU activation and subsequently fine-tune the entire model with a learning rate of $5\times 10^{-5}$. For QuALITY, we concatenate the document vector with an explicit encoding of the question and candidate answer and predict four scores via softmax on the minibatch-level, one per question-answer pair. For MuLD-Mov., we instead concatenate the encoded question: "Is the character <NAME> behaving like a hero or a villain?", where <NAME> is replaced by the actual character’s name. We use a sigmoid function (threshold at 0.5) to predict the binary label.
@@ 3.3 Baselines
(Nussbaum et al., 2024) is a text embedding model for document lengths up to 8,192 tokens. Based on a BERT architecture, the model is trained via Masked Language Modeling and both unsupervised as well as supervised contrastive textual similarity tasks. We choose this model as a highly competitive baseline with a long context window yet relatively small size at 137M parameters.
@@ 4 Results
- MiniLM-256 384 65.45 85.57 83.49 85.70 31.23
- MPNet-256 768 05.97 11.03 58.14 73.53 27.74
- DistilRoBERTa-256 768 43.05 66.75 58.14 73.20 29.51
SBERT: MiniLM 384 30.87 48.35 58.14 73.53 31.96
SBERT: MPNet 768 44.26 67.17 58.14 73.53 33.05
SBERT: DistilRoBERTa 768 27.95 45.91 58.14 73.53 32.83
Table 3: Evaluation results (in %, averaged over five fine-tuning runs) for NextLevelBERT pretrained on the text chunk embeddings of various encoder models. All next-level models use text chunks of size 256.
@@ 4.1 Downstream Task Evaluation
On the MuLD-Mov. dataset, five out of seven NextLevel models as well as Nomic Embed are able to surpass chance level (58%, i.e., a majority class classifier). All other models fail to do so, even the Avg. SBERT baseline, which does well on the other datasets. Notably, our sentence-based and 256er-chunked NextLevelBERT models also outperform the Longformer baseline from Hudson and Al Moubayed (2022) (82.58 F1), the original benchmark paper, without any task-specific preprocessing. To attain their score, the authors first extracted the most important passages based on character name frequency. We hypothesize that a good performance on MuLD is dependent on a model’s ability to process the entire text as well as exchange contextual information internally.
On the QuALITY dataset, Nomic Embed is the strongest model, presumably due to its sufficient context size for this dataset and the fact that it can provide token-level details to answer the questions. The Avg. SBERT baselines mostly reach a higher accuracy than NextLevelBERT. Since the chunks that are mean-pooled by Avg. SBERT are also the inputs to NextLevelBERT, a lower performance suggests that NextLevelBERT may actively disregard some of the incoming information. This is likely due to the fact that during pretraining, NextLevelBERT learned to focus on reconstructing general events and to ignore local details that would be hard to predict from the chunk-level context. QuALITY consists of questions such as “Why does Joseph lie about the water supply?”, which require non-trivial consideration of local details. In general, this is a difficult dataset for all the tested models and the results are generally barely above the chance level of 25%. This is partly reflected in the dataset’s leaderboard [^4] https://nyu-mll.github.io/quality/, where end-to-end encoder-only models without additional preprocessing or fine-tuning on related datasets report similar accuracy. The dataset’s current SOTA mostly relies on large ($\geq$7B) decoder-only models, training on a large-scale task-specific reading comprehension corpus, and/or retrieval augmentation.
For the same effective context size of 8,192 tokens, Nomic Embed outperforms NextLevelBERT-16 on all datasets. This and the overall results indicate that NextLevelBERT makes heavy use of the pretrained encoder model’s ability to contextualize the incoming chunk’s text. Smaller chunk sizes, such as 16, likely limit this effectiveness.
@@ 4.2 Impact of Chunking Strategy
The results in Table 2 illustrate how the optimal choice for dividing the input text into chunks can vary between use cases. We hypothesize that there is a trade-off between increasing the receptive field by using larger chunks and losing information due to higher levels of compression necessary to encode longer text chunks in fixed-size sentence embeddings. This is supported by the fact that in the cases where the next-level models’ context size is not sufficient, performance quickly degrades. A notable exception is NextLevelBERT with sentence-based chunking on the MuLD dataset. With an average context size of $\sim$12k tokens and MuLD’s average document length being 34,706 tokens, it still outperforms almost all other models. A possible explanation is that the movie scripts that MuLD consists of are dialogue-heavy and the sentence-based chunking plus the contextualization abilities of the model contribute to successfully predicting the role of the character without seeing the whole document.
BookSum MuLD-Mov QuALITY
NextLeveBERT MRR HR@10 Acc F1 Acc
with 15% 65.52 85.57 80.93 83.45 31.88
with 30% 61.14 82.10 69.07 75.86 30.08
with 40% 64.21 84.81 72.79 77.54 29.53
simpler masking 00.77 01.29 58.14 73.53 24.69
random init. 01.18 01.78 58.14 73.45 26.44
Table 4: Results for alternative masking rates (1-3), replacing all selected inputs with the [MASK] vector (4), and a randomly initialized model (5). All models use text chunks of size 256. We report the average over five fine-tuning runs.
@@ 4.3 Impact of Encoder Model
Downstream results for NextLevelBERT with different text encoder models are compared in Table 3. We observe that the effectiveness of the original encoder model does not automatically translate to similar outcomes in the next-level setting. Surprisingly, pretraining on MiniLM leads to the best downstream results, although it is the smallest model in terms of parameters as well as embedding dimensionality. A possible explanation is that capturing the intricacies of a 384-dimensional embedding space during pretraining is easier than those of a 768-dimensional space.
@@ 4.4 Impact of Document Length on Model Performance
To investigate how the models perform with regard to the length of the document, we visualize the MRR on BookSum per document length in Figure 2. Here, 2(a) provides a general plot covering all document lengths, while 2(b) shows a more fine-grained view of documents shorter than 768 tokens to study if the design for long documents comes at the cost of degraded performance on shorter texts. NextLevelBERT outperforms the other models, except Nomic Embed, across all document lengths with a margin of up to 20% MRR. Even on the shorter documents, it generally improves over the other models starting from 341 tokens. On the longest documents, all models exhibit a MRR of 1, but this is due to the low number of samples at these lengths. Notably, Longformer seems unable to generate any useful document embeddings for any sequence length out-of-the-box.
@@ 4.5 Pretraining Effects
To measure the effect of initializing NextLevelBERT with the chunk encoder weights, we train two versions of the same model, one initialized with the weights of the sentence encoder and one with random initialization. We compare their downstream performance in Table 4. For the MuLD-Movie and the QuALITY tasks, the model with random initialization does not score above chance level. On BookSum it also underperforms, indicating that it was not able to learn meaningful representations. Together, these results reveal that initializing NextLevelBERT with the chunk encoder weights is essential to its downstream capabilities.
@@ 5 Related Work
5.1 Efficient Attention
Numerous approaches have been proposed to increase the context length by making the attention mechanism more scalable, typically by sparsifying or approximating the attention operation: Beltagy et al. (2020) use a combination of local and global attention to circumvent the quadratic scaling. They train the Longformer, a BERT-based LM that can incorporate a context size of up to 4,096 tokens, showing improved results on long document classification. Building on the Longformer, Xiao et al. (2022) pack multiple documents into the same sequence and mask salient sentences to generate multi-document summaries. The Poolingformer (Zhang et al., 2021) adds a second level of pooling-combined attention to achieve linear computational complexity, while the Reformer (Kitaev et al., 2019) proposes attention with locality-sensitive hashing to reduce the attention operation to linearithmic time. Finally, HyperAttention (Han et al., 2023) efficiently approximates the attention matrix and enables computing attention in near-linear time.
5.2 Hierarchical Approaches for Long Contexts
@@
5.3 Embedding-based Loss Objectives
The process of encoding sequences of data and predicting the next item in the sequence’s latent representation has also been used in other domains. To generate images autoregressively, Lee et al. (2022) use a variation of a Vector-Quantized Variational Autoencoder (VAE) to encode sequences of image patches into discretized codebook representations. Their model learns to predict the latent representation of the next patch in the sequence. The decoder of the VAE then outputs pixel-level images from the generated latent patch sequence. In reinforcement learning, Lee et al. (2022) use cubic masking patches to corrupt video sequences as an auxiliary task for learning a state representation model. The model learns to reconstruct the uncorrupted videos’ latent representations. Similar approaches have also been suggested in Hafner et al. (2020, 2022). In NLP, sentence embedding approaches, such as DeCLUTR (Giorgi et al., 2021) and Denoising Autoencoder with Transformers (Moradshahi et al., 2023), learn to generate text embeddings for semantic similarity via a self-supervised loss on the embedding space. DeCLUTR samples text pairs that are adjacent, overlapping, or where one sample subsumes the other to obtain positive pairs for a contrastive loss. The Denoising Autoencoder corrupts input sentences and learns to reconstruct the original sentence.
@@ 6 Conclusion
Successfully modeling long-range dependencies is an important step towards building language models with general language understanding. Inspired by the idea of modeling language at a coarser resolution than that of tokens, we have proposed NextLevelBERT, and transferred masked language modeling from the token level to general text representation vectors. We pretrained NextLevelBERT on a large collection of books and evaluated the resulting text representations on three diverse downstream tasks in a long document setting. Our results show that next-level masked language modeling leads to sensibly contextualized text chunk representations and generates useful long-document vectors, especially for tasks that do not require specific token-level information. Additionally, NextLevelBERT is very parameter-efficient, using only a third of the parameters of the Nomic Embed model while maintaining similar performance.
In the future, combining the predictions of a next-level model with a token-level component based on the required granularity could be a way of preserving token-level details. Furthermore, pretraining NextLevelBERT on data from more diverse domains could improve its performance in general and extend its usefulness to a wider range of tasks. Lastly, while not straightforward, adapting decoder-only models to the next level could translate similar gains as in the MLM setup to generative language models.